at cruising speed, and airplane flew 500 kilometers with the wind in the same amount of time it took to fly 400 kilometers against the wind. if the constant speed of the wing is 20kph, what is the cruising speed of the airplane?
Answer by mananth(16949)

at cruising speed, and airplane flew 500 kilometers with the wind in the same amount of time it took to fly 400 kilometers against the wind. if the constant speed of the wind is 20kph, what is the cruising speed of the airplane?
let speed be x
With wind speed = x+20
time = 500 / x+20
against wind it will be x-20
time = 400 /x-20
500 / x+20 = 400 / x-20
500(x-20)=400(x+20)
500x-10000 = 400x+8000
100x = 18000
x= 180 mph